# Competition for binding targets results in paradoxical effects for simultaneous activator and repressor action - Extended Version

### v1

@@ III-B Dynamics flip during initial transient
To examine the onset of the paradoxical effect, we determine the small-time relationship between the output $Y$ and the input $u_{A}$. We can represent the system under consideration as $\dot{x}=f(x,p)$ where $p$ is the set of parameters (i.e., rate constants) that the system depends on and $x=\{T_{R},T_{A},D_{R},D_{A},Y,A_{{\textsc{t}}},R_{{\textsc{t}}}\}$. We define $Y(x(t))$ as the projection onto the $Y$ species. We indicate the Lie derivative by $L_{{\scriptscriptstyle f(x,p)}}Y(x(t))=(\nabla Y)f(x,p)=\frac{dY}{dt}$.
@@
$L_{{\scriptscriptstyle f(x,p)}}(L_{{\scriptscriptstyle f(x,p)}}Y(x(t)))=\\ \beta(-(t_{r}^{+}(T_{{\textsc{t}}}-T_{R}-T_{A})(R_{{\textsc{t}}}-D_{R}-T_{R})-t_{r}^{-}T_{R}-\bar{\delta}T_{R})\\ -(t_{a}^{+}(T_{{\textsc{t}}}-T_{R}-T_{A})(A_{{\textsc{t}}}-D_{A}-T_{A})-t_{a}^{-}T_{A}-\bar{\delta}T_{A}))\\ +\kappa(t_{a}^{+}(T_{{\textsc{t}}}-T_{R}-T_{A})(A_{{\textsc{t}}}-D_{A}-T_{A})-t_{a}^{-}T_{A}-\bar{\delta}T_{A})\\ -\gamma(\beta(T_{{\textsc{t}}}-T_{R}-T_{A})+\kappa T_{A}-\gamma Y)$ (17)
@@
Figure 8: Transient dynamics is devoid of paradoxical effects. Time series of the (a) concentration of the output protein $Y(t)$ for $u_{A}=0.1$ and $u_{A}=100$, and (b) difference between the output levels of the protein $Y(t,u_{A}=0.1)-Y(t,u_{A}=100)$, for $K_{ta}=50$. Generally, one may describe the existence for initial crossing times as depicted in (c). The parameter values are same as Figure 7
Figure 8 displays the flip in the behavior of the synthetic circuit after an initial transient to exhibit the paradoxical effect. The time series of $Y(t)$ shows the existence of the paradoxical effect, where the output level for $u_{A}=0.1$ is higher than that of $u_{A}=100$. However, taking a closer look at the data, the effect emerges after a finite duration of time. This is shown by plotting the difference between the output levels for $u_{A}=0.1$ and $u_{A}=100$ in Figure 8(b). We see that $t<t_{o}$ (where $t_{o}>0$ is a threshold value of time for different combinations of inputs, $t_{0}\approx 0.15$ for $u_{A}=0.1$ and $u_{A}=100$), $Y(u_{A}=100)>Y(u_{A}=0.1)$, i.e., an increase in the activator increases the output protein levels and therefore no paradoxical effect. Whereas for $t>t_{o}$, $Y(u_{A}=100)<Y(u_{A}=0.1)$, i.e., an increase in the activator decreases the output protein levels and therefore exhibits the paradoxical effect. Hence, the output protein levels of the circuit with low input values (say $u_{A}=0.1$) start at $t=0$ with a lower initial growth rate, cross the trajectory of an intermediate input level (say $u_{A}=100$), before reaching a higher steady state value than the intermediate one, as depicted in Figure 8(c). The parameters used for the numerical simulations are somewhat arbitrary, therefore the magnitude of the effects and the timescales could be expected to be different in practice.
@@ III-C Increasing decoy sites increases the output
The proposed synthetic circuit exhibits the paradoxical effect given by:
[…]
where $\bar{Y}$ is the steady-state levels of the protein when other inputs ($u_{R},u_{A}$) and parameters ($T_{{\textsc{t}}}$, reaction rate constants) are kept constant.
Proof.  Under the limit of $K_{ta}\rightarrow\infty,\overline{T_{A}}\rightarrow\infty$, and from equation (6), the steady-state levels of the output protein is:
@@
$\overline{Y}=\frac{\beta\overline{T}}{\gamma}\implies\frac{\mathrm{d}\overline{Y}}{\mathrm{d}D_{{\textsc{t}}}}=\frac{\beta}{\gamma}\frac{\mathrm{d}\overline{T}}{\mathrm{d}D_{{\textsc{t}}}}.$
The conservation at steady state shown in equations (11)-(13) becomes:
$\overline{A_{{\textsc{t}}}}$ $=\overline{A}+\frac{\overline{D}\overline{A}}{K_{da}},$ $\overline{R_{{\textsc{t}}}}=$ $\overline{R}+\frac{\overline{T}\overline{R}}{K_{tr}}+\frac{\overline{D}\overline{R}}{K_{dr}},$ (19)
$\overline{T_{{\textsc{t}}}}$ $=\overline{T}+\frac{\overline{T}\overline{R}}{K_{tr}},$ $\overline{D_{{\textsc{t}}}}=$ $\overline{D}+\frac{\overline{D}\overline{A}}{K_{da}}+\frac{\overline{D}\overline{R}}{K_{dr}},$ (20)
$\therefore\overline{A}$ $=\frac{\overline{A_{{\textsc{t}}}}}{1+\frac{\overline{D}}{K_{da}}},$ $\overline{R}=$ $\frac{\overline{R_{{\textsc{t}}}}}{1+\frac{\overline{D}}{K_{dr}}+\frac{\overline{T}}{K_{tr}}}.$ (21)
Substituting equations (21) in (20), and differentiating with respect to $D_{{\textsc{t}}}$, we get:
$C_{1}\frac{\mathrm{d}\overline{D}}{\mathrm{d}D_{{\textsc{t}}}}=C_{2}+C_{3}\frac{\mathrm{d}\overline{T}}{\mathrm{d}D_{{\textsc{t}}}}\quad\text{and}\quad C_{4}\frac{\mathrm{d}\overline{T}}{\mathrm{d}D_{{\textsc{t}}}}=C_{5}\frac{\mathrm{d}\overline{D}}{\mathrm{d}D_{{\textsc{t}}}},$
where
$C_{1}$ $=1-\frac{D_{{\textsc{t}}}}{\left(1+\frac{\overline{R}}{K_{dr}}+\frac{\overline{A}}{K_{da}}\right)^{2}}\left[\frac{\overline{R}^{2}}{K_{dr}^{2}R_{{\textsc{t}}}}+\frac{\overline{A}^{2}}{K_{da}^{2}A_{{\textsc{t}}}}\right]>0,$
$C_{2}$ $=\frac{1}{1+\frac{\overline{R}}{K_{dr}}+\frac{\overline{A}}{K_{da}}}>0,$
$C_{3}$ $=\frac{D_{{\textsc{t}}}}{\left(1+\frac{\overline{R}}{K_{dr}}+\frac{\overline{A}}{K_{da}}\right)^{2}}\frac{\overline{R}^{2}}{K_{dr}K_{tr}R_{{\textsc{t}}}}>0,$
$C_{4}$ $=1-\frac{\overline{T}^{2}\overline{R}^{2}}{K_{tr}^{2}T_{{\textsc{t}}}R_{{\textsc{t}}}}>0\quad\text{and}\quad C_{5}=\frac{\overline{T}^{2}\overline{R}^{2}}{K_{tr}K_{dr}T_{{\textsc{t}}}R_{{\textsc{t}}}}>0.$
Eliminating $\frac{\mathrm{d}\overline{D}}{\mathrm{d}D_{{\textsc{t}}}}$, we get:
$\frac{\mathrm{d}\overline{T}}{\mathrm{d}D_{{\textsc{t}}}}=\frac{C_{2}}{\left(\frac{C_{1}C_{4}}{C_{5}}-C_{3}\right)}>0,\quad\text{as}\quad\frac{C_{1}C_{4}}{C_{5}}-C_{3}>0.$
Therefore, under the limit of $K_{ta}\rightarrow\infty$,
@@
$\frac{\mathrm{d}\overline{D}}{\mathrm{d}D_{{\textsc{t}}}}>0.$
@@
Figure 7 (b) shows the variation of steady-state output levels ($\overline{Y}$) as a function of decoy sites ($D_{{\textsc{t}}}$) for different values of $K_{ta}$. For $K_{ta}=0.1$, we observe the expected behavior of the output levels decreasing as the amount of decoy sites increases. On the contrary, the behavior flips for higher $K_{ta}$ values showing the exhibition of the paradoxical effect. In the two-parameter bifurcation plots in Figure 7 (c), we see that for high $K_{ta}$ value, the paradoxical affect is observed for all values of $u_{A}$. Increasing the amount of $D_{\textsc{t}}$, increases the output concentration. However, the value of $D_{\textsc{t}}$ after which the system moves to an activated state depends on the value of $u_{A}$. Therefore, the onset of the effect can be controlled using the activator levels. In the case of low $K_{ta}$ value in Figure 7 (d), we see that increasing $D_{{\textsc{t}}}$, decreases the protein levels irrespective of the value of $u_{A}$.
@@
$\partial_{D_{{\textsc{t}}}}Y(0,D_{{\textsc{t}}})=\partial_{D_{{\textsc{t}}}}\dot{Y}(0,D_{{\textsc{t}}})=\partial_{D_{{\textsc{t}}}}\ddot{Y}(0,D_{{\textsc{t}}})=0,$
@@
$D_{{\textsc{t}}}(-(\kappa-\beta)t_{a}^{+}(T_{{\textsc{t}}}-T_{R}-T_{A})d_{a}^{+}(A_{{\textsc{t}}}-D_{A}-T_{A})\,+\\ \beta t_{r}^{+}(T_{{\textsc{t}}}-T_{R}-T_{A})d_{r}^{+}(R_{{\textsc{t}}}-D_{R}-T_{R}))\\ =D_{{\textsc{t}}}(-(\kappa-\beta)t_{a}^{+}d_{a}^{+}TA+\beta t_{r}^{+}d_{r}^{+}TR)$ (22)
@@ IV Additional system properties
Proof.  From Claim 1, there is a compact and convex set of values of our system species $\{T_{R},T_{A},D_{R},D_{A},Y,A_{{\textsc{t}}},R_{{\textsc{t}}}\}$ that is closed under the time evolution of our system. By Brouwer’s fixed point theorem for every $t\geq 0$ we have the time evolution $\phi_{t}$ has a fixed point $\bar{x}_{t}$. Take a sequence of $t_{n}\rightarrow 0$, the sequence $\bar{x}_{t_{n}}$ has a limit point $\bar{x}$. If $\bar{x}$ was not an equilibrium, then for small $t$ it would not be a fixed point for $\phi_{t}$, which is a contradiction. Thus $\bar{x}$ is an equilibrium of our vector field.
@@
Proof.  Using the conservation laws, $T_{{\textsc{t}}}=T+T_{A}+T_{R}$ and $D_{{\textsc{t}}}=D+D_{A}+D_{R}$, we can form the “extended rate function” by simply replacing the dynamics for $\dot{T}$ and $\dot{D}$ with our conservation laws. We can then take the determinant of the Jacobian of this modified rate function and verify all its terms are positive, which implies injectivity by Theorem 8.1 in [12]. Mathematica code for this computation can be found in the appendix. Inspection of the output confirms the Jacobian is always nonzero and thus our system is injective.

### v2

@@ III-B Dynamics flip during initial transient
One might wonder at what time can one observe the paradoxical effect. To examine the onset of the paradoxical effect, we determine the small-time relationship between the output $Y$ and the input $u_{A}$. We can represent the system under consideration as $\dot{x}=f(x,p)$ where $p$ is the set of parameters (i.e., rate constants) that the system depends on and $x=\{T_{R},T_{A},D_{R},D_{A},Y,A_{{\textsc{t}}},R_{{\textsc{t}}}\}$. We define $Y(x(t))$ as the projection onto the $Y$ species. We indicate the Lie derivative by $L_{{\scriptscriptstyle f(x,p)}}Y(x(t))=(\nabla Y)f(x,p)=\frac{dY}{dt}$.
@@
$L_{{\scriptscriptstyle f(x,p)}}(L_{{\scriptscriptstyle f(x,p)}}Y(x(t)))\\ =\beta(-(t_{r}^{+}(T_{{\textsc{t}}}-T_{R}-T_{A})(R_{{\textsc{t}}}-D_{R}-T_{R})-t_{r}^{-}T_{R}-\bar{\delta}T_{R})-(t_{a}^{+}(T_{{\textsc{t}}}-T_{R}-T_{A})(A_{{\textsc{t}}}-D_{A}-T_{A})-t_{a}^{-}T_{A}-\bar{\delta}T_{A}))\\ +\kappa(t_{a}^{+}(T_{{\textsc{t}}}-T_{R}-T_{A})(A_{{\textsc{t}}}-D_{A}-T_{A})-t_{a}^{-}T_{A}-\bar{\delta}T_{A})-\gamma(\beta(T_{{\textsc{t}}}-T_{R}-T_{A})+\kappa T_{A}-\gamma Y)$ (22)
@@
Figure 7: Transient dynamics is devoid of paradoxical effects. Time series of the (a) concentration of the output protein $Y(t)$ for $u_{A}=0.1$ and $u_{A}=100$, and (b) difference between the output levels of the protein $Y(t,u_{A}=0.1)-Y(t,u_{A}=100)$, for $K_{ta}=50$. Generally, one may describe the existence for initial crossing times as depicted in (c). The parameter values are same as Figure 6.
Figure 7 displays the flip in the behavior of the synthetic circuit after an initial transient to exhibit the paradoxical effect. The time series of $Y(t)$ shows the existence of the paradoxical effect, where the output level for $u_{A}=0.1$ is higher than that of $u_{A}=100$. However, taking a closer look at the data, the effect emerges after a finite duration of time. This is shown by plotting the difference between the output levels for $u_{A}=0.1$ and $u_{A}=100$ in Figure 7-b. We see that $t<t_{o}$ (where $t_{o}>0$ is a threshold value of time for different combinations of inputs, $t_{0}\approx 0.15$ for $u_{A}=0.1$ and $u_{A}=100$), $Y(u_{A}=100)>Y(u_{A}=0.1)$, i.e., an increase in the activator increases the output protein levels and therefore no paradoxical effect. Whereas for $t>t_{o}$, $Y(u_{A}=100)<Y(u_{A}=0.1)$, i.e., an increase in the activator decreases the output protein levels and therefore exhibits the paradoxical effect. Hence, the output protein levels of the circuit with low input values (say $u_{A}=0.1$) start at $t=0$ with a lower initial growth rate, cross the trajectory of an intermediate input level (say $u_{A}=100$), before reaching a higher steady state value than the intermediate one, as depicted in Figure 7-c. The parameters used for the numerical simulations are somewhat arbitrary, therefore the magnitude of the effects and the timescales could be expected to be different in practice.
@@ III-C Increasing decoy sites increases the output
The output $\overline{Y}$ of our synthetic circuit depends on $D_{\textsc{t}}$ as follows:
$\frac{\mathrm{d}\overline{Y}}{\mathrm{d}D_{\textsc{t}}}=C_{3}+\frac{C_{4}}{K_{ta}}$ (24)
where,
$C_{3}$ $=\frac{\frac{C_{5}}{K_{tr}\gamma}(\frac{\kappa\overline{A}}{K_{ta}}+\beta)\frac{T_{\textsc{t}}}{O^{2}_{t}}}{1+\frac{D_{\textsc{t}}}{K_{dr}O^{2}_{d}}(1+\frac{\overline{A}}{K_{da}})+\frac{T_{\textsc{t}}}{K_{tr}O^{2}_{t}}(1+\frac{\overline{A}}{K_{ta}})}$
@@
$C_{4}$ $=\frac{\mathrm{d}\overline{A}}{\mathrm{d}D_{\textsc{t}}}\frac{T_{\textsc{t}}}{\gamma O^{2}_{t}}\left[\kappa-\beta+\frac{\kappa\overline{R}}{K_{tr}}\right]$
$C_{5}$ $=\frac{\overline{R}}{K_{dr}O_{d}}-\left[\frac{\overline{R}D_{\textsc{t}}}{K_{dr}K_{da}O^{2}_{d}}+\frac{\overline{R}D_{\textsc{t}}}{K_{tr}K_{ta}O^{2}_{t}}\right]\frac{\mathrm{d}\overline{A}}{\mathrm{d}D_{\textsc{t}}}$
@@
with $\frac{\mathrm{d}\overline{A}}{\mathrm{d}D_{\textsc{t}}}<0$ defined in Appendix -B and $\overline{A}$ and $\overline{R}$ depend on $D_{\textsc{t}}$.
Remark 2
Notice that $C_{3}>0$ and $C_{4}<0$. This means that $\frac{\mathrm{d}\overline{Y}}{\mathrm{d}D_{\textsc{t}}}$ is the sum of a positive term and a negative term. In particular, for small values of $K_{ta}$ the negative second term dominates, while for large values of $K_{ta}$ the positive first term dominates, such that:
$\frac{\mathrm{d}\overline{Y}}{\mathrm{d}D_{\textsc{t}}}>0\quad\text{for}\quad k_{ta}\rightarrow\infty,$
contradicting the expectation that increasing the competitor sites ($D_{\textsc{t}}$) decreases the output $\overline{Y}$.
Proof.  Differentiating equation (14) with respect to $D_{\textsc{t}}$ we have
$\frac{\mathrm{d}\overline{Y}}{\mathrm{d}D_{\textsc{t}}}=\frac{1}{\gamma}\left[\frac{\kappa\overline{A}}{K_{ta}}+\beta\right]\frac{\mathrm{d}\overline{T}}{\mathrm{d}D_{\textsc{t}}}+\frac{1}{\gamma}\frac{\kappa\overline{T}}{K_{ta}}\frac{\mathrm{d}\overline{A}}{\mathrm{d}D_{\textsc{t}}}$
Using a similar approach to Theorem 1, we compute $\frac{\mathrm{d}\overline{A}}{\mathrm{d}D_{\textsc{t}}}$ and $\frac{\mathrm{d}\overline{T}}{\mathrm{d}D_{\textsc{t}}}$ to get our output in the form (24).
@@
Figure 6-b shows the variation of steady-state output levels ($\overline{Y}$) as a function of decoy sites ($D_{{\textsc{t}}}$) for different values of $K_{ta}$. For $K_{ta}=0.1$, we observe the expected behavior of the output levels decreasing as the amount of decoy sites increases. On the contrary, the behavior flips for higher $K_{ta}$ values showing the exhibition of the paradoxical effect. In the two-parameter bifurcation plots in Figure 6-c, we see that for high $K_{ta}$ value, the paradoxical affect is observed for all values of $u_{A}$. Increasing the amount of $D_{\textsc{t}}$, increases the output concentration. However, the value of $D_{\textsc{t}}$ after which the system moves to an activated state depends on the value of $u_{A}$. Therefore, the onset of the effect can be controlled using the activator levels. In the case of low $K_{ta}$ value in Figure 6-d, we see that increasing $D_{{\textsc{t}}}$, decreases the protein levels irrespective of the value of $u_{A}$.
@@
$\partial_{D_{{\textsc{t}}}}Y(0,D_{{\textsc{t}}})=\partial_{D_{{\textsc{t}}}}\dot{Y}(0,D_{{\textsc{t}}})=\partial_{D_{{\textsc{t}}}}\ddot{Y}(0,D_{{\textsc{t}}})=0.$
@@
$D_{{\textsc{t}}}(-(\kappa-\beta)t_{a}^{+}(T_{{\textsc{t}}}-T_{R}-T_{A})d_{a}^{+}(A_{{\textsc{t}}}-D_{A}-T_{A})\,+\beta t_{r}^{+}(T_{{\textsc{t}}}-T_{R}-T_{A})d_{r}^{+}(R_{{\textsc{t}}}-D_{R}-T_{R}))\\ =D_{{\textsc{t}}}(-(\kappa-\beta)t_{a}^{+}d_{a}^{+}TA+\beta t_{r}^{+}d_{r}^{+}TR)$ (25)
@@ IV Additional system properties
The following claims give some additional information about the qualitative behavior of the system. In particular we can rigorously prove that the system has bounded trajectories, and that there exists a unique equilibrium.
@@
Proof.  From Claim 1, there is a compact and convex set of values of our system species $\{T_{R},T_{A},D_{R},D_{A},Y,A_{{\textsc{t}}},R_{{\textsc{t}}}\}$ that is closed under the time evolution of our system. By Brouwer’s fixed point theorem for every $t\geq 0$ we have the time evolution $\phi_{t}$ has a fixed point $\overline{x}_{t}$. Take a sequence of $t_{n}\rightarrow 0$, the sequence $\overline{x}_{t_{n}}$ has a limit point $\overline{x}$. If $\overline{x}$ was not an equilibrium, then for small $t$ it would not be a fixed point for $\phi_{t}$, which is a contradiction. Thus $\overline{x}$ is an equilibrium of our vector field.
@@
Proof.  Using the conservation laws, $T_{{\textsc{t}}}=T+T_{A}+T_{R}$ and $D_{{\textsc{t}}}=D+D_{A}+D_{R}$, we can form the “extended rate function” by simply replacing the dynamics for $\dot{T}$ and $\dot{D}$ with our conservation laws. We can then take the determinant of the Jacobian of this modified rate function and verify all its terms are positive, which implies injectivity by Theorem 8.1 in [12]. Mathematica code for this computation can be found in [13]. Inspection of the output confirms the Jacobian is always nonzero and thus our system is injective.